An uncharged capacitor and a resistor are connected in series to a source of EMF. If ε = 7.38 V, C = 21.1 μF, and R = 152 Ω, calculate the time constant τ of the circuit.

Calculate the maximum charge on the capacitor.

Calculate the charge on the capacitor after one time constant.
Answer #1

Time constant=C*R=21.1E-6*152=0.0032

Maximum charge on capacitor=21.1E-6*152=0.0032 coulomb

Voltage after one constant=7.38*(1-e-1)=4.665 V

Charge after chargong=21.1e-5*4.665=98.4e-6 coulomb